OpenStudy (anonymous):

what is the resultant of the given vectors as an ordered pair? < -7,9 > and <-1,-6>

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[r=(-7i+9j)+(-1i-6j)=(-7-1)i+(9-6)j=-8i+3j\] its magnitude \[\left| vector~ r \right|=\left| \sqrt{ \left( -8 \right)^2+3^2} \right|=\sqrt{73} \]

OpenStudy (anonymous):

is u saying 73 the answer @surjithayer

OpenStudy (anonymous):

it is magnitude. r vector is <-8,3>